# Autonomous Exploration and Semantic Updating of Large-Scale Indoor Environments with Mobile Robots

### v1

@@ IV-A Quantitative Analysis
Construction Update I Update II
Classes GT Map GT Map GT Map
Tables 45 45 45 39 45 37
Chairs 109 84 99 73 109 69
Doors 210 153 210 155 210 156
TABLE I: Semantic map object count across the construction and update stages. GT: Ground Truth, Map: Mapping result. The table compares the ground truth with the actual results across the construction, update I, and update II scenarios.
The robot traverses the environment, updating the semantic map based on the observed changes. The results of the semantic map construction and updates are presented in Table I. We can see that the detection of chairs and doors is relatively more challenging. During update I, the removal of 10 chairs after the construction phase is reflected in the decrease in the number of chairs from 84 to 73. However, in Update II, only 69 chairs were detected, differing from the expected result due to occlusions. Table II presents the breakdown of object detection performance across the three experimental phases. Doors have a relatively low mAP, mainly due to the it’s partial views while robot is traversing in narrow corridors. For chairs, the accuracy decreases during the update stages, indicating the challenges of occlusions.
Table Chair Door mAP
Scenario TP FP FN TP FP FN TP FP FN Table Chair Door
Construction 36 9 4 84 0 16 140 13 68 0.90 0.84 0.67
Update I 36 3 9 72 1 23 140 15 66 0.80 0.76 0.68
Update II 33 4 6 67 2 29 142 14 68 0.85 0.69 0.68
TABLE II: Object detection performance and mAP analysis. TP: True Positives, FP: False Positives, FN: False Negatives, mAP: Mean Average Precision for detected objects across the construction and update stages.
@@ IV-B Qualitative Analysis
Figure 6: Examples of accurate detection
@@
Fig. 6 illustrates several examples of accurate object detection during our experiments for chairs, tables, and doors. In Fig. 7, we discuss representative challenging scenarios in object detection during robot navigation.
Occlusion Challenges. Objects like tables and chairs being in close proximity to each other, often results in occlusion. It leads to objects not being detected or segmented as a single object. This can be seen in Fig. 7(a) where table and chairs are segmented as table. In Fig. 7(c) table, chair, and a portion of the pillar are segmented together as table. In Fig. 7(b), a chair is partially occluded by the pillar. This can lead to incorrect object associations when the chair is detected from another view. In Fig. 7(e), the chairs in the background are completely occluded by the table and are not observed by the robot. Subsequently they are not associated with the semantic map.
Limitations in Constrained Spaces. The presence of constrained spaces like narrow corridors in the environment can hinder the robot’s ability to view the complete structure of objects such as doors as seen in Fig. 7(f). Many such doors are not detected, leading to reduction in the number of object nodes in the semantic map. This is evident in Table I.
Deviation from the Traversal Trajectory. In Fig. 7(d), large bin in the corridor occludes the objects behind it from the robot view, forces the robot to take a detour. This detour results in some objects being seen by the robot. Such scenarios greatly impact the quality of the semantic mapping, reducing the ability to accommodate environment changes.
Impact of Lighting. Varying lighting conditions can cause shadows and reflections. As seen in Fig. 7(g) and Fig. 7(h), these shadows on the floor are mistook for a door and a table, respectively, reducing the overall detection and semantic map accuracy. In Fig. 7(c), the overall environment brightness is reduced during the night, contributing to poor detection and segmentation.
Structural Misidentification. Walls resemble doors when viewed at certain angles, having similar rectangular structure. Some walls also have additional structures like support bars or cable cord covers, resembling large door handles. Hence, these walls are often misidentified as doors. One such case can be seen in Fig. 7(i).
@@ V Conclusion and Future Work
We introduce a robotic system that enables a mobile robot to autonomously explore an unknown environment, recognize objects in the environment, and subsequently build a semantic map of the environment. The system utilizes a hybrid representation of the semantic map that consists of an occupancy grid map for geometry and a topological map for semantics. This representation enables us to update the semantic map efficiently to reflect the changes in the environment. Experiments conducted in a large indoor environment demonstrate the effectiveness of our system in autonomous exploration and semantic mapping.
Object perception during robot navigation still remains a challenge. Our future work will consider interactive object recognition, in which a robot can plan its actions to detect objects such as looking at some objects, moving towards an object to confirm its existence, and moving around to avoid occlusions.
Acknowledgement

### v2

@@ IV-A Quantitative Analysis
Table II summarizes the number of objects detected in both Environment A and Environment B during the Construction, Update I and Update II phases. In Environment A, we can see that while doors and tables were stably detected, the detection of chairs was relatively more challenging. During update I, the removal of 10 chairs after the construction phase is reflected in the decrease in the number of chairs detected from 84 to 73. However, in Update II, only 69 chairs were detected, with false negatives increasing to 29 as shown in Table III. This was mainly because some relocated chairs were placed in the path of the robot traversal trajectory. As a result, robot took a detour to avoid hitting them, causing it to miss some chairs that were not relocated and previously detected. Moreover, people occupying chairs occluded them, increasing false negatives. In contrast, Environment B has very stable detections, with many classes having 100% precision. However, some categories like cabinets that are dark colored, are challenging to detect in nominal lightning, resulting in them not being detected.
@@ IV-B Qualitative Analysis
Figure 8: Examples of accurate detection and segmentation of chairs, tables and doors in Environment A.
Class Construction Update I Update II
TP FP FN Precision TP FP FN Precision TP FP FN Precision
Environment A
Table 36 9 4 0.9 36 3 9 0.8 33 4 6 0.85
Chair 84 0 16 0.84 72 1 23 0.76 67 2 29 0.69
Door 140 13 28 0.83 140 15 26 0.84 142 14 28 0.84
Environment B
Person 2 0 0 1 2 0 0 1 2 0 0 1
Trash Bin 2 0 0 1 0 0 0 - 0 1 0 0
Umbrella 1 0 0 1 0 0 0 - 0 0 0 -
Chair 3 0 0 1 3 0 0 1 1 0 0 1
Television 1 1 0 0.5 1 0 0 1 1 0 0 1
Cabinet 2 0 1 0.67 1 0 2 0.33 2 0 1 0.67
TABLE III: Semantic map object count across the construction and update stages for A & B Environments. TP: True Positives, FP: False Positives, FN: False Negatives, Precision: Precision for detected objects across the construction and update stages.
In Fig. 9, we show several challenging scenarios in object detection and segmentation, while robot is navigating.
Occlusion Challenges. Close proximity of objects such as tables and chairs causes occlusion and leads to missed detections or incorrect segmentation, as seen in Fig. 9(a). In Fig. 9(c) table, chair, and a portion of the pillar are segmented together as table. The scenario seen in Fig. 9(b), where a chair is partially occluded by the pillar, leads to incorrect object associations when the chair is detected again at another view.
Detection Threshold. Lowering the detection threshold increases false positives, while a higher threshold may miss objects. In Fig. 9(e), chairs in the foreground are undetected at a 0.8 threshold. careful choice of threshold for object class is important for accurate semantic mapping.
Limitations in Constrained Spaces. Narrow corridors limit the robot’s view of objects like doors, as seen in Fig. 9(f). Several of such doors are not detected, leading to a reduction in the number of object nodes in the semantic map. This is evident in Table II.
@@
Deviation from the Traversal Trajectory. In Fig. 9(d), a large bin in the corridor occluded objects, forcing the robot to take a detour. This detour resulted in the robot not seeing some objects. Such scenarios impact the quality of the semantic mapping, reducing the ability to accommodate environmental changes.
Impact of Lighting. Varying lighting conditions can cause shadows and reflections. As seen in Fig. 9(g) and Fig. 9(h), these shadows on the floor were mistook for a door and a table, respectively, reducing the overall detection and semantic map accuracy. In Fig. 9(c), reduced brightness at night degraded detection and segmentation.
Structural Misidentification. Walls resembling doors due to shape or added features (e.g., support bars) often cause misidentification, as seen in Fig. 9(i). This scenario also occurred in Environment B, update II phase, where a cardboard box was identified as a trash bin, leading to a false positive.
@@ V Conclusion and Future Work
We introduce a robotic system that enables a mobile robot to autonomously explore an unknown environment, recognize objects in the environment, and subsequently build a semantic map of the environment. The system utilizes a hybrid representation of the semantic map that consists of an occupancy grid map for geometry and a topological map for semantics. This representation enables us to update the semantic map efficiently to reflect the changes in the environment. Experiments conducted on a mobile manipulatory system in a large- and medium-scale indoor environments demonstrate the effectiveness of our system in autonomous exploration and semantic mapping.
Object perception during robot navigation still remains a challenge. Our future work will consider interactive object recognition, in which a robot can plan its actions to detect objects such as looking at some objects, moving towards an object to confirm its existence for a robust object association, and dynamic planning of traversal to avoid occlusions.
Acknowledgment